OpenStudy (anonymous):

A park in a subdivision is triangular-shaped. Two adjacent sides of the park are 573 feet and 536 feet. The angle between the sides is 58mc023-1.jpg. To the nearest unit, find the area of the park in square yards. A) 32,557 yd2 B) 14,470 yd2 C) 28,940 yd2 D) 43,410 yd2

OpenStudy (asadkarim7):

i dont know the angle yo u have posted can u tell me the magnitude of the angle

OpenStudy (anonymous):

it is a 58 degree angle

OpenStudy (anonymous):

area = (product of sides times sine of the included angle)/2 = (A*B sinC)/2 where A, B are the sides and C is the angle

OpenStudy (asadkarim7):

then you should apply the formula 1/2 a b sin 58 and convert the answer from feet to yards ^2

OpenStudy (anonymous):

okay I think I can try to figure that out

OpenStudy (asadkarim7):

you want me to solve that for you >??

OpenStudy (anonymous):

if you could, these are the ones i'm stuck on

OpenStudy (asadkarim7):

ok wait let me bring my calculator

OpenStudy (asadkarim7):

1/2 (573)(536) sin 58= 130229.66

OpenStudy (asadkarim7):

convert 130229.66 to yard^2

OpenStudy (anonymous):

you have to divide that by 9 to get the answer in square yards.

OpenStudy (anonymous):

i'd say B

OpenStudy (anonymous):

great thanks!

OpenStudy (asadkarim7):

np (: